Fermi National Laboratory
 

Accelerator Update for May 24, 2006 - May 25, 2006

 

Wednesday May 24

The day shift began with Booster turn on underway and with techs accessing the Tevatron (TeV) to work on a few limited jobs.

Booster experts began a study period at 8:10 AM.

The Main Injector (MI) began taking beam and sent some on to the Recycler at 8:37 AM.

Water technicians turned on the MINOS absorber room pumps at 9 AM.

At 9:40 AM, Operations began preparing MI and Booster for accesses. MI techs will check on the 603 resistor and Booster techs will conduct quadrupole shunt investigations.

Booster techs came out of access at 10:48 AM. Operations restored beam about ten minutes later.

The Collider Detector at Fermilab (CDF) powered their solenoids at 11:46 AM.

Operations established beam to MI at 1:11 PM. Experts began studies.

Operations powered all the TeV bulk power supplies at 8:18 PM.

Thursday May 25

The midnight shift began with Operations monitoring a small Recycler stash of protons, and with MI and Recycler experts conducting studies.

MI studies ended at 12:51 AM.

TeV techs reported at 2:18 AM that they completed testing all the HFUs and they all had tested good.

Operations had the Booster, MI, and TeV ready for accesses by 6:07 AM.

The Plans for Today

The plans are to allow MI experts to access their tunnels. The Run Coordinator will allow an access tomorrow to fix the P2 vacuum leak.

Machine Reports

Linac

Everything is running normally. They conducted a steering study. The Neutron Therapy Facility will start treating patients today at 10 AM.

Booster

The machine is running well. Techs will begin disconnecting shunts on ORBMP today.

MI

Experts have moved the tunes back to where they should be. Two quadrupole magnets have to be moved to correct the orbit in the vertical plane.

Pbar

Experts will run mock Recycler transfers today.

TeV

Experts hope to power the TeV tomorrow. Vacuum techs are pumping down the B11 separator. Experts are considering putting the tank through another bake.

Recycler

Experts are ready for more protons.

Switchyard 120

Experts will access the P3 line this morning and searched for the soft ground fault.

MiniBooNE

Experts are ready for beam.

NuMI

Experts report that their safety system tests are finished. They plan to conduct power supply tests next Tuesday, 5/30/06.

Collider Detector at Fermilab

Experts reported that they had to pull a plug. The work should be completed and the plug moved back in by next Wednesday, 5/31/06.

D-Zero

Their detector has been closed.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/25/2006    email Fermilab